OpenStudy (compassionate):

One of the first electronic computers was in the shape of a huge box. It was 96 feet long and 2.5 feet wide. The amount of space inside was approximately 3,000 cubic feet. How many feet high was the computer?

OpenStudy (compassionate):

So 2.5 * 96 * x = 3,000 3000/240= 12.5 Can someone explain why?

OpenStudy (mayankdevnani):

see amount of space inside means VOLUME

OpenStudy (mayankdevnani):

in your question,Volume is given and two sides are given and then find other side

OpenStudy (compassionate):

Hmm... I kinda get it.

OpenStudy (mayankdevnani):

REMEMBER :- \[\large \bf VOLUME=length \times breadth \times height\]

OpenStudy (mayankdevnani):

plug all the values in it and get your answer :)

OpenStudy (jonnyvonny):

>breadth >mfw

OpenStudy (mayankdevnani):

breadth means WIDTH

OpenStudy (compassionate):

>mfw meme arrows

OpenStudy (compassionate):

Righto. I get it. Danke

OpenStudy (mayankdevnani):

good :)

OpenStudy (triciaal):